OpenStudy (anonymous):

If the game only has 20 levels, explain how to find the value of the series for the coins and the points. Use complete sentences and arrive at final values. I already did the coins one by myself im having trouble with this part because i am plugging my data into a1–a1(r)^(n) / 1- r. My formula (3-3(3)^(3))/(1-3) for the third level. Level Points 1 3 2 9 3 27

OpenStudy (anonymous):

but i keep getting 39 instead of 27 @phi

OpenStudy (chmvijay):

@Ahunter2222 @cwrw238 @ganeshie8 @djcool31 @e.cociuba @fallonbanks @samigupta8 can u help him :)

OpenStudy (phi):

I posted in the chem section. But to repeat, your formula finds the sum of the points 3+9+27= 39 you want the nth value, which is 3^n

OpenStudy (anonymous):

yes level 20 is the n value

OpenStudy (anonymous):

im a retard

OpenStudy (anonymous):

wait nvm I need to solve for the third level so I can show this formula works

OpenStudy (anonymous):

I am not getting 27 when i input 3 as the n value showing that my formula does not work

OpenStudy (phi):

First, what are you computing? if you want to list the number of points you get at level n, it will be 3^n that formula says that at level 3, you will get 3^3 = 27 points on the other hand, if you want to find the total number of points you get from level 1 to level 3, you get 3+9+27= 39 points

OpenStudy (anonymous):

ok now I am a retard so my equation was right all along i just forgot that i was adding them.

OpenStudy (phi):

It's impolite to call people names. But if you get confused, it is helpful to think about how you got confused... what did you forget or mis-read... and try to fix that thinking so next time you know what to do.

OpenStudy (anonymous):

Well thank you for the help

OpenStudy (anonymous):

wait can you tell me if my restrictions are correct?

OpenStudy (phi):

what are the restrictions ?

OpenStudy (anonymous):

The restrictions on domain of the coins table would be there cannot be any negative levels. Restrictions on the range would be no odd or negative numbers. For the second table the restrictions for the domain would be only positive integers.

OpenStudy (anonymous):

Level Coins 1 2 2 4 3 6 Level Points 1 3 2 9 3 27 these are both of my tables

OpenStudy (anonymous):

actually positive integers for both the domain and range on the second

OpenStudy (phi):

yes, they sound good. Though the range for the coins is positive even numbers, {2,4,6...} it might be ok to just say positive integers

OpenStudy (anonymous):

okay thanks

OpenStudy (phi):

yw